OpenStudy (anonymous):

f(x,y) = (5/2)e^(x^2+y^2) dydx from y = -sqrt(4-x^2) to y = sqrt(4-x^2) from x = - 2 to x = 2

OpenStudy (anonymous):

I need help integrating.

OpenStudy (anonymous):

I know the answer, confirmed by Wolfram Alpha and the solutions guide, I just don't know how to get there.

OpenStudy (anonymous):

Have you converted to polar?

OpenStudy (anonymous):

Cylindrical coordinates make it easier but I would like to keep it in cartesian.

OpenStudy (anonymous):

then: \(5/2\int_{-\sqrt(4-x^2)}^{\sqrt(4-x^2)}\int_{-2}^{2}e^{x^2}e^{y^2}dxdy\)

OpenStudy (anonymous):

without an extra y or x, how do you differentiate e^(x^2) or e^(y^2)?

OpenStudy (anonymous):

A change in coordinates will simplify the work. \[x=r\cos\theta\\ y=r\sin\theta\\ x^2+y^2=r^2\\ dy~dx=r~dr~d\theta\] Then change the limits accordingly.

OpenStudy (anonymous):

Thanks Siths but the question wants to see both versions. It's goal is to demonstrate how changing coordinate systems simplifies things.

OpenStudy (anonymous):

Well @myko's integral gives you the Cartesian integral. Apply the conversions and you'll get your polar integral.

OpenStudy (anonymous):

They can't really expect you to evaluate the Cartesian one, can they?

OpenStudy (anonymous):

you will not find it easy to integrate in cartesian, Look about erf(x) function in wikipedia

OpenStudy (anonymous):

I'll ask the instructor because we never learned the error function.

OpenStudy (anonymous):

\[\frac{5}{2}\int_{-\sqrt{4-x^2}}^\sqrt{4-x^2}\int_{-2}^2e^{x^2+y^2}~dy~dx=\frac{5}{2}\color{red}{\int_{0}^{2\pi}}\color{blue}{\int_{0}^{2}}e^{r^2}~r~\color{blue}{dr}~\color{red}{d\theta}\]

OpenStudy (anonymous):

Should be \(dx~dy\), but you get the idea...

OpenStudy (anonymous):

I think you are missing an r to account for the Jacobian transformation but I got it. Thanks.

OpenStudy (anonymous):

No, it's there. \(dx~dy=\color{red}{r}~dr~d\theta\).